# Proving the existence of localized patterns and saddle node bifurcations in 1D activator-inhibitor type models

### v1

@@ 5.2 Bounds for Saddle Nodes
In this section, we compute the bounds of Theorem 5.1 in order to prove existence and uniqueness of zeros to (7). As a result, we obtain a rigorous enclosure of a saddle node bifurcation to (2). We will begin with the $\mathcal{Y}_{0}$ bound. This can be done by combining the approach in Section 6 of unbounded_domain_cadiot to treat a scalar PDE with an extra equation and that of Section 4.1 of gs_cadiot_blanco where a system of PDEs was considered.
@@
The proof follows from Lemma 6.5 of unbounded_domain_cadiot . In particular, each $\mathcal{Z}_{u,k,j}$ can be computed using the results of the aforementioned lemma. ∎
@@ 5.3 Proofs of Saddle Nodes
Choose $N_{0}=800,N=620,d=40$. Then, we perform the full construction to build $\overline{\mathbf{x}}=\overline{\text{{$\gamma$}}}^{\dagger}(\nu,\overline{\mathbf{U}},\overline{\mathbf{W}})$. Then, we define $\mathbf{x}_{glys}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\mathbf{x}}$. Next, we construct $B^{N}$ and use Lemma 2.2 to compute $\kappa$. This allows us to compute the $\mathcal{Z}_{2}(r)$ bound defined in Section 5.2. Finally, using julia_blanco_fassler , we choose $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4\times 10^{-8}$ and define
@@ 6.1 Control of the eigenvalues
Now in order to obtain a precise control on the set $)\sigma_{\delta_{0}}\cap\sigma(D\mathbb{f}(\tilde{\mathbf{u}}$)), we use the spectrum of $Df(\bar{\mathbf{U}})$, as presented in cadiot2025stabilityanalysislocalizedsolutions . For this purpose, we first compute a pseudo-diagonalization of $Df(\bar{\mathbf{U}})$, that is $P$ and $P^{-1}$ such that
@@
Then, using similar notations as in cadiot2025stabilityanalysislocalizedsolutions , we define $S$ as the diagonal matrix having the same diagonal as $\mathcal{D}$, that is
@@
The bounds $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3}$ introduced in the previous lemma are not the same as the ones introduced in Lemma 4.7. However, they are closely related and can be computed in a very similar fashion. Indeed, their computation is based on the proof of Lemma 4.7, combined with the estimates of Lemma 6.4. Note that the additional factor $2$ in the inequalities for $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3}$ allows to provide upper bounds for $\mathcal{Z}_{u,1}^{(q)},\mathcal{Z}_{u,2}^{(q)},\mathcal{Z}_{u,3}^{(q)}$ given in Lemma 4.2 of cadiot2025stabilityanalysislocalizedsolutions (see the proof of Theorem 5.2 in the aforementioned paper). Details for the computation of the bounds of Lemma 6.5 are given in cadiot2025stabilityanalysislocalizedsolutions , and are implemented explicitly in the code julia_blanco_fassler . In fact, such computations heavily relies on the analysis presented in Section 4, and we prefer not to present them in details to avoid intensive overlapping.
@@ 6.3 Stability analysis
In this section, we focus on the stability of some localized stationary solutions of Section 3. In particular, given $\tilde{\mathbf{u}}$ proven in Section 3, that is there exists $r_{0}$ such that $\|\tilde{\mathbf{u}}-\bar{\mathbf{u}}\|_{\mathcal{H}}\leq r_{0}$, then we use Lemma 6.5 to control the spectrum of $D\mathbb{f}(\tilde{\mathbb{u}}):L^{2}_{e}\to L^{2}_{e}$. In fact, this allows us to conclude about the (nonlinear) stability of $\tilde{\mathbf{u}}$ with respect to even perturbations in $L^{2}_{e}$. In the cases for which $\tilde{\mathbf{u}}$ is unstable, we count the number of unstable directions. More specifically, we obtain the following results for which computational details are given in julia_blanco_fassler .
@@ 7 Conclusion
In this manuscript, we further investigated the results of the authors of AlSaadi_unifying_framework . More specifically, the authors of the aforementioned work computed solutions to (2) on long 1D domains along with various bifurcations. The authors asked if their results could be rigorously justified, and we provided the justification for some of their results. Indeed, we provided proofs of existence and local uniqueness of some of their approximate solutions on the unbounded domain $\mathbb{R}$. Moreover, we provide the general framework needed to attempt to prove any of the approximate localized patterns in AlSaadi_unifying_framework . Following this, we proved the existence of a saddle-node bifurcation on the unbounded domain $\mathbb{R}$. This not only provides an answer to one type of bifurcation studied in AlSaadi_unifying_framework , but also provides the framework needed to prove saddle-node bifurcations on unbounded domains using the theory of unbounded_domain_cadiot . Our results have provided partial answers to some of the questions asked by the authors of AlSaadi_unifying_framework as well as an additional application of the method developed in unbounded_domain_cadiot to saddle-node bifurcations.
@@
While we have answered some of the questions asked by the authors of AlSaadi_unifying_framework , there are many more still present. In particular, the authors computed various branches of solutions to (2) by performing numerical continuation. Some of these branches lead to saddle-node bifurcations, which could now be proven using our approach described in Section 5; however, this does not allow for the proof of the branch itself. The necessary theory to obtain proofs of branches of solutions on unbounded domains has already been developed and applied in whitham_cadiot . In particular, the approach of Section 4.6 in whitham_cadiot allowed for one to obtain the proof of the branch in the Whitham equation. This approach relies on the method demonstrated in breden_polynomial_chaos and marschal . The approach would need to be modified slightly as whitham_cadiot used parameter continuation. This approach works well assuming no bifurcations are encountered, which is not the case here as we have saddle-node bifurcations. Combining the approach used in marschal along with the method for unbounded domains used in whitham_cadiot would allow for a rigorous proof of the branch. We consider this a potential future work.
@@
Another question asked by the authors of AlSaadi_unifying_framework was regarding Hopf bifurcations. In particular, the authors describe the location of various approximate Hopf bifurcations in many of the models that fit into (2). Furthermore, they ask about whether or not these Hopf bifurcations are sub/super critical. The approach to prove Hopf bifurcations on unbounded domains would be similar to that for saddle-node bifrucations except with a more complicated augmented system. This work is in-progress.
@@
To further the last point, associated to each Hopf bifurcation is a branch of time-periodic solutions. This branch was not computed by the authors of AlSaadi_unifying_framework . We consider it of interest to investigate such a branch of solutions. Furthermore, could we obtain a rigorous proof of a solution along this branch? In the framework of computer assisted proofs, this would be similar to proving a solution that is localized in its spatial dimensions but periodic in the time dimension. Proving such a solution has not been considered using the approach of unbounded_domain_cadiot . This work is also in-progress.
@@
To conclude, while our work here has provided a partial answer to some of the questions asked by the authors of AlSaadi_unifying_framework , we believe many more of their results could be rigorously verified using similar methods to those described in this manuscript. We believe our methods could be of use to those studying models of similar type to (2). For instance, proofs for the 2D Gray-Scott model were already provided in gs_cadiot_blanco . We believe this manuscript opens many possibilities for rigorous investigation of these models.
@@ 8 Acknowledgments
The authors wish to thank Professor Nicolas Verschueren van Rees for his insights and fruitful discussions.
@@ Appendix A Computing the Bound $\mathcal{Z}_{2}$ for Saddle Nodes
where $\mathcal{Z}_{2,2}(r)$ is defined as in (139). Notice that due to our choice of estimate for $\|z_{i,j}h_{k,l}h_{m,n}\|_{2},$ we obtain the same result for each quantity. This significantly simplifies our computation compared to the approach used in gs_cadiot_blanco . It also allow us to use the same bound for the saddle node bounds as we did for patterns. This is one of the reasons we chose this approach in this work.
@@ Appendix B Integral Computation using residues
These formulas are then implemented and evaluated using interval arithmetic in julia_blanco_fassler .

### v2

@@ 5.2 Bounds for Saddle Nodes
In this section, we compute the bounds of Theorem 5.1 in order to prove existence and uniqueness of zeros to (7). As a result, we obtain a rigorous enclosure of a saddle node bifurcation to (2). We will begin with the $\mathcal{Y}_{0}$ bound. This can be done by combining the approach in Section 6 of Cadiot et al. (2024) to treat a scalar PDE with an extra equation and that of Section 4.1 of Cadiot and Blanco (2025) where a system of PDEs was considered.
@@
The proof follows from Lemma 6.5 of Cadiot et al. (2024). In particular, each $\mathcal{Z}_{u,k,j}$ can be computed using the results of the aforementioned lemma. ∎
@@ 5.3 Proofs of Saddle Nodes
Choose $N_{0}=800,N=620,d=40$. Then, we perform the full construction to build $\overline{\mathbf{x}}=\overline{\text{{$\gamma$}}}^{\dagger}(\nu,\overline{\mathbf{U}},\overline{\mathbf{W}})$. Then, we define $\mathbf{x}_{glys}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\mathbf{x}}$. Next, we construct $B^{N}$ and use Lemma 2.2 to compute $\kappa$. This allows us to compute the $\mathcal{Z}_{2}(r)$ bound defined in Section 5.2. Finally, using Blanco et al. (2025), we choose $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4\times 10^{-8}$ and define
@@ 6.1 Control of the eigenvalues
Now in order to obtain a precise control on the set $)\sigma_{\delta_{0}}\cap\sigma(D\mathbb{f}(\tilde{\mathbf{u}}$)), we use the spectrum of $Df(\bar{\mathbf{U}})$, as presented in Cadiot (2025b). For this purpose, we first compute a pseudo-diagonalization of $Df(\bar{\mathbf{U}})$, that is $P$ and $P^{-1}$ such that
@@
Then, using similar notations as in Cadiot (2025b), we define $S$ as the diagonal matrix having the same diagonal as $\mathcal{D}$, that is
@@
The bounds $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3}$ introduced in the previous lemma are not the same as the ones introduced in Lemma 4.7. However, they are closely related and can be computed in a very similar fashion. Indeed, their computation is based on the proof of Lemma 4.7, combined with the estimates of Lemma 6.4. Note that the additional factor $2$ in the inequalities for $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3}$ allows to provide upper bounds for $\mathcal{Z}_{u,1}^{(q)},\mathcal{Z}_{u,2}^{(q)},\mathcal{Z}_{u,3}^{(q)}$ given in Lemma 4.2 of Cadiot (2025b) (see the proof of Theorem 5.2 in the aforementioned paper). Details for the computation of the bounds of Lemma 6.5 are given in Cadiot (2025b), and are implemented explicitly in the code Blanco et al. (2025). In fact, such computations heavily relies on the analysis presented in Section 4, and we prefer not to present them in details to avoid intensive overlapping.
@@ 6.3 Stability analysis
In this section, we focus on the stability of some localized stationary solutions of Section 3. In particular, given $\tilde{\mathbf{u}}$ proven in Section 3, that is there exists $r_{0}$ such that $\|\tilde{\mathbf{u}}-\bar{\mathbf{u}}\|_{\mathcal{H}}\leq r_{0}$, then we use Lemma 6.5 to control the spectrum of $D\mathbb{f}(\tilde{\mathbb{u}}):L^{2}_{e}\to L^{2}_{e}$. In fact, this allows us to conclude about the (nonlinear) stability of $\tilde{\mathbf{u}}$ with respect to even perturbations in $L^{2}_{e}$. In the cases for which $\tilde{\mathbf{u}}$ is unstable, we count the number of unstable directions. More specifically, we obtain the following results for which computational details are given in Blanco et al. (2025).
@@ 7 Conclusion
In this manuscript, we further investigated the results of the authors of Al Saadi et al. (2021). More specifically, the authors of the aforementioned work computed solutions to (2) on long 1D domains along with various bifurcations. The authors asked if their results could be rigorously justified, and we provided the justification for some of their results. Indeed, we provided proofs of existence and local uniqueness of some of their approximate solutions on the unbounded domain $\mathbb{R}$. Moreover, we provide the general framework needed to attempt to prove any of the approximate localized patterns in Al Saadi et al. (2021). Following this, we proved the existence of a saddle-node bifurcation on the unbounded domain $\mathbb{R}$. This not only provides an answer to one type of bifurcation studied in Al Saadi et al. (2021), but also provides the framework needed to prove saddle-node bifurcations on unbounded domains using the theory of Cadiot et al. (2024). Our results have provided partial answers to some of the questions asked by the authors of Al Saadi et al. (2021) as well as an additional application of the method developed in Cadiot et al. (2024) to saddle-node bifurcations.
@@
While we have answered some of the questions asked by the authors of Al Saadi et al. (2021), there are many more still present. In particular, the authors computed various branches of solutions to (2) by performing numerical continuation. Some of these branches lead to saddle-node bifurcations, which could now be proven using our approach described in Section 5; however, this does not allow for the proof of the branch itself. The necessary theory to obtain proofs of branches of solutions on unbounded domains has already been developed and applied in Cadiot (2025a). In particular, the approach of Section 4.6 in Cadiot (2025a) allowed for one to obtain the proof of the branch in the Whitham equation. This approach relies on the method demonstrated in Breden (2023) and Calleja et al. (2024). The approach would need to be modified slightly as Cadiot (2025a) used parameter continuation. This approach works well assuming no bifurcations are encountered, which is not the case here as we have saddle-node bifurcations. Combining the approach used in Calleja et al. (2024) along with the method for unbounded domains used in Cadiot (2025a) would allow for a rigorous proof of the branch. We consider this a potential future work.
@@
Another question asked by the authors of Al Saadi et al. (2021) was regarding Hopf bifurcations. In particular, the authors describe the location of various approximate Hopf bifurcations in many of the models that fit into (2). Furthermore, they ask about whether or not these Hopf bifurcations are sub/super critical. The approach to prove Hopf bifurcations on unbounded domains would be similar to that for saddle-node bifrucations except with a more complicated augmented system. This work is in-progress.
@@
To further the last point, associated to each Hopf bifurcation is a branch of time-periodic solutions. This branch was not computed by the authors of Al Saadi et al. (2021). We consider it of interest to investigate such a branch of solutions. Furthermore, could we obtain a rigorous proof of a solution along this branch? In the framework of computer assisted proofs, this would be similar to proving a solution that is localized in its spatial dimensions but periodic in the time dimension. Proving such a solution has not been considered using the approach of Cadiot et al. (2024). This work is also in-progress.
@@
To conclude, while our work here has provided a partial answer to some of the questions asked by the authors of Al Saadi et al. (2021), we believe many more of their results could be rigorously verified using similar methods to those described in this manuscript. We believe our methods could be of use to those studying models of similar type to (2). For instance, proofs for the 2D Gray-Scott model were already provided in Cadiot and Blanco (2025). We believe this manuscript opens many possibilities for rigorous investigation of these models.
@@ 8 Acknowledgments
The authors wish to thank Professor Nicolas Verschueren van Rees for his insights and fruitful discussions. MC was supported by the ANR project CAPPS: ANR-23-CE40-0004-01 and by the FMJH : ANR-22-EXES-0013.
@@ Appendix A Computing the Bound $\mathcal{Z}_{2}$ for Saddle Nodes
where $\mathcal{Z}_{2,2}(r)$ is defined as in (139). Notice that due to our choice of estimate for $\|z_{i,j}h_{k,l}h_{m,n}\|_{2},$ we obtain the same result for each quantity. This significantly simplifies our computation compared to the approach used in Cadiot and Blanco (2025). It also allow us to use the same bound for the saddle node bounds as we did for patterns. This is one of the reasons we chose this approach in this work.
@@ Appendix B Integral Computation using residues
These formulas are then implemented and evaluated using interval arithmetic in Blanco et al. (2025).